OpenStudy (anonymous):

Find the population standard deviation (σ not S) for the data set: {75,80,90,95,75,85,55,80,90,95} Round your answer to the nearest 10th.

OpenStudy (welshfella):

first find the mean of these numbers can you do that?

OpenStudy (anonymous):

no

OpenStudy (welshfella):

add them all up and divide by 10

OpenStudy (anonymous):

82 is what I got

OpenStudy (welshfella):

yes you are right

OpenStudy (welshfella):

now you need to find the difference between each number in the list and 82 so for first number 75 it is 82 - 75 = 7 for 80 it is 82 - 80 = 2 do this for all the other numbers

OpenStudy (welshfella):

90 - 82 = 8 and so on

OpenStudy (anonymous):

75=7 80=2 90=8 95=13 75=7 85=3 55=27 80=2 90=8 95=13

OpenStudy (welshfella):

right strictly it should have been 80 - 82 = -2 but its not that important because we square each of those results

OpenStudy (welshfella):

so now we square ach of the 10 results 7^2 = 49 2^2 = 4 and so on

OpenStudy (welshfella):

when you have done that add them all up then divide the answer by the number of values (10) this result is called the variance find the square root of the variance to get the standard deviation

OpenStudy (welshfella):

use your calculator - it would be a bit long winded otherwise. It a bit long winded anyway!

OpenStudy (welshfella):

in the formula x is a number , mu is the mean and N is the total numbers.

OpenStudy (anonymous):

75=7^2=49 80=2^2=4 90=8^2=64 95=13^2=169 75=7^2=49 85=3^2=9 55=27^2=729 80=2^2=4 90=8^2=64 95=13^2=169 1310 divided by 10=131

OpenStudy (welshfella):

the biig 'E' means 'the sum of'.

OpenStudy (welshfella):

right i got that too so std dev = sqrt 131

OpenStudy (anonymous):

I'm confused on how to do the equation

OpenStudy (welshfella):

mu is the mean x - mu is the differences oops sorry i made a mistake with that - the formula is just a summary of the process you have to go through the procedure as we did

OpenStudy (welshfella):

correct is